OpenStudy (anonymous):

The formula A = 118e0.024t models the population of a particular city, in thousands, t years after 1998. When will the population of the city reach 140 thousand? A. 2008 B. 2005 C. 2006 D. 2007 @ganeshie8

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

Hey

OpenStudy (anonymous):

set \(118e^{.024t}=140\) and solve for \(t\) it takes three steps

OpenStudy (anonymous):

first step it to divide by \(118\) to get \[e^{.024t}=\frac{140}{118}=\frac{70}{59}\]

OpenStudy (anonymous):

second step it to write in equivalent exponential form as \[.024t=\ln(\frac{70}{59})\]

OpenStudy (anonymous):

then divide the result by \(.024\) to get \[t=\frac{\ln(\frac{70}{49})}{.024}\] i guess the next step is to use a calculator to find that number

OpenStudy (anonymous):

looks like it is about \(7\) http://www.wolframalpha.com/input/?i=ln%2870%2F59%29%2F.024

OpenStudy (whpalmer4):

except the last bit should be \[t = \frac{\ln(\frac{70}{59})}{0.024}\]not \(49\) in the denominator as a finger slip put it in the previous post.

OpenStudy (anonymous):

and 7 years after 1998 is ... oh typo sorry good catch @whpalmer4

OpenStudy (whpalmer4):

Makes a big difference in the result if you use 49!

OpenStudy (anonymous):

i am sure

OpenStudy (anonymous):

Would the answer be B?

OpenStudy (anonymous):

is 2005 a choice?

OpenStudy (anonymous):

Yes. That's answer B.

OpenStudy (anonymous):

ok pick that one hope the steps were more or less clear

OpenStudy (anonymous):

They were. Thank you!